OpenStudy (bethgrace96):

which expression below is equivalent to 4i(12-7i) a)-28+48i b)28+48i c)48+28i d)-48=28i

jimthompson5910 (jim_thompson5910):

When you distribute 4i(12-7i), what do you get?

OpenStudy (bethgrace96):

48i-28i^2?

jimthompson5910 (jim_thompson5910):

what is i^2 equal to

OpenStudy (bethgrace96):

i+i? lol idk

jimthompson5910 (jim_thompson5910):

i by defintion is equal to sqrt(-1) in other words, \[\Large i = \sqrt{-1}\]

OpenStudy (bethgrace96):

oh wait its 1

jimthompson5910 (jim_thompson5910):

so square both sides to get \[\Large i = \sqrt{-1}\] \[\Large i^2 = (\sqrt{-1})^2\] \[\Large i^2 = -1\]

OpenStudy (bethgrace96):

so the answer is c?

jimthompson5910 (jim_thompson5910):

Since \[\Large i^2 = -1\] we go from 48i-28i^2 to 48i-28(-1)

jimthompson5910 (jim_thompson5910):

now simplify 48i-28(-1)

OpenStudy (bethgrace96):

oh okay so its b?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (bethgrace96):

thanks a bunch!!!

jimthompson5910 (jim_thompson5910):

you're welcome